OpenStudy (anonymous):

(Modeling with Quadratic Equations Help) A rocket is fired upward. At the end of the burn it has an upward velocity of 147 m/sec and is 73.5 meters high. Find it it's maximum height when it is attained.

OpenStudy (anonymous):

So here is my work so far... f(x)=ax^2+bx+c s=-4.9t^2+VoT + h I plugged in the numbers for the equation s=-4.9t^2+147t+73.5 I began to solve to complete the square.. s=-4.9(t^2+147t)+73.5 s-73.5=-4.9(t^2+147t) What do i do from here?

OpenStudy (anonymous):

oops meters

OpenStudy (acxbox22):

you dont need to complete the square..you can just divide the whole equation by -4.9 then factor then solve

OpenStudy (anonymous):

forget completing the square first coordinate of the vertex is always \(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case \[-\frac{147}{2\times (-4.9)}\] whatever that is

OpenStudy (anonymous):

The whole section we are studying in advanced algebra is on completing the square. The example in the book, like this question, completes the square. @acxbox22

OpenStudy (anonymous):

I finished completing the square and got an answer of 1176 meters in 15 seconds. I googled the question and another person got same answer. Though it does not make since since the maximum height was 73.5 meters high in 147 seconds so in 15 seconds, that number should be much lower than 73.5